OpenStudy (anonymous):

11^x=12 In terms of logarithms

OpenStudy (anonymous):

\[b^x=y\iff x=\log_b(y)\]

OpenStudy (anonymous):

you should be able to switch back and forth quickly

OpenStudy (anonymous):

by "should" i mean it would help you immensely in your math class if you can go back and forth without agony

OpenStudy (anonymous):

yeah still work in progress haha just started learning them

OpenStudy (anonymous):

i keep getting the incorrect answer I am unsure of how to enter it into the calculator. DO i have to type 11 before the log(12)?

OpenStudy (anonymous):

it is not a calculator exercise unless you are solving for x

OpenStudy (anonymous):

Iforgot to mention that I am solving for x, sorry.

OpenStudy (anonymous):

oh i misread. you are solving for x right?

OpenStudy (anonymous):

ok then \[b^x=y=\iff x=\frac{\ln(y)}{\ln(b)}\]

OpenStudy (anonymous):

change 'o base formula

OpenStudy (anonymous):

\[x=\frac{\ln(12)}{\ln(11)}\]

OpenStudy (anonymous):

ohhhhh i see! how did you know to change it to natural log/natural log?

OpenStudy (anonymous):

You could have used \[\log_{10} \]You get the same answer. \[x=\log_{10}12/\log_{10}11 \]